# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 2.3. Differences with Existing Work
Previous surveys of LLMs for recommender system usually categorized existing work with a classification standard and introduced studies respectively. Lin et al.  (Lin et al., 2023a) categorized existing work into the targets and the methods of adapting LLMs to recommendation tasks. Wu et al.  (Wu et al., 2023b) mainly focused on the form of information between LLMs and recommender systems, including LLM embeddings, LLM tokens and LLMs. Fan et al.  (Fan et al., 2023) summarized the framework into four sections, involving deep representation learning, pre-training, fine-tuning and prompting of LLMs. These research mainly concentrated on demonstrating related work and summarizing the advantages and limitations.
Compared to previous work, we concentrate on the ability of LLMs leveraged for recommendation tasks, and provide a systematic empirical analysis on prompting LLM-based recommendations by devising a general framework ProLLM4Rec. We mainly focus on two aspects, i.e., LLMs and prompt engineering, providing definitions and solutions from both conceptual and methodological perspectives. Furthermore, we conduct experiments to discover new findings and validate results previously discussed in existing research, serving as an inspiration for future research efforts.
@@ 3. GENERAL FRAMEWORK AND OVERALL SETTINGS
Figure 1. The overall framework of our proposed ProLLM4Rec.
In this section, we present our proposed ProLLM4Rec with two important components, namely LLMs and prompts. Generally speaking, the rich world knowledge and general capabilities of LLMs demonstrate the potential to develop LLM-powered recommender systems. Nevertheless, it is essential to introduce appropriate prompting engineering to provide domain-specific knowledge on recommendation tasks for LLMs to serve as personalized recommenders. In the following, we first describe our general framework by defining several key elements (Section 3.1). Then, we explain how our framework can be generalized to various recommendation scenarios and tasks by framework instantiation (Section 3.2). Finally, we introduce the details of the overall experimental settings for further analysis (Section 3.3).
@@ 3.1. Overview of the Approach
When it comes to the use of LLMs in recommender systems, the first step involves choosing suitable LLMs tailored to specific scenarios, with an emphasis on their distinct capabilities. Subsequently, it is important to conduct prompt engineering to instruct LLMs to perform effective recommendations. Typically, in the realm of recommender systems using LLMs, despite the task settings in existing studies vary a lot, the overall prompting format remains relatively consistent with minor variations (Hou et al., 2023; Gao et al., 2023; Dai et al., 2023a; Wang et al., 2023c). Therefore, to unify existing prompting approaches for different recommendation purposes, as illustrated in Fig. 1, we establish a general framework for prompt engineering in ProLLM4Rec, consisting of four key factors in the prompts. Firstly, task description is required to clearly express the specific goal of recommendation tasks. Secondly, prompts need to be carefully designed, to express user interest and meanwhile enable LLMs to provide personalized recommendations with both world and domain knowledge. Thirdly, the purpose of a recommender system is to provide users with candidate items, and potential candidate items should be constructed for LLMs to facilitate the recommendation with the understanding of domain-specific item information. Furthermore, special prompting strategies can be further employed to enhance the specialized recommendation capabilities of LLMs. In what follows, we will thoroughly examine the impact of each factor on the recommendation performance.
@@ 3.1.1. Key Elements in Our Framework
To carry out the experiments, we first describe the key elements in ProLLM4Rec, to clarify their definitions and scope in this work. Specially, we introduce the following five elements for our framework:
@@
$\bullet$ Large language models. As proposed in previous research (Hou et al., 2023; Zhang et al., 2023b; Li et al., 2023a; Bao et al., 2023a), there exists a large gap between general language modeling and personalized user behavioral modeling, making it non-trivial to utilize LLMs in recommender systems. In this work, we investigate the efficacy of LLMs from perspectives of public availability, tuning strategies, model architecture, parameter scale, and context length, aiming to gain insights into the selection of appropriate LLMs for performing recommendation tasks. Observations and discussions on the use of LLMs are presented in Section 4.
@@
$\bullet$ Task description. To adapt LLMs to the scenario of recommendation, it is necessary to clearly express the context and target of recommender systems for LLMs in the prompt, i.e., task description. With different prompt descriptions of tasks, large language models can be utilized to various recommendation scenarios and tasks such as click-through rate predictions (Kang et al., 2023; Bao et al., 2023b, a), sequential recommendation (Hou et al., 2023; Dai et al., 2023a; Zheng et al., 2023) and conversational recommender systems (He et al., 2023; Gao et al., 2023; Wang et al., 2023h).
@@
$\bullet$ User interest modeling. The modeling of user interest is the key to recommendation tasks (Rendle et al., 2009; He et al., 2017). When leveraging LLMs for recommendation, users are generally expressed in natural language text, which is different from traditional approaches capturing user preference from ID-based behavior sequences (Kang and McAuley, 2018; Wu et al., 2021a). In this paper, we mainly consider the reflected user interest based on his or her interaction behaviors with interacted items. Especially, as detailed in Section 5.2, we employ item description texts, user profiles, and historical interactions between users and items to reveal the underlying user interest using natural languages (Wang et al., 2023c; Lin et al., 2023b; Shu et al., 2023; Yao et al., 2023).
@@
$\bullet$ Candidate items construction. The purpose of recommender systems is to provide users with items to choose from, so candidate items construction is a crucial step in our framework (Hou et al., 2023; Zhang et al., 2023f; Dai et al., 2023a). A simple approach is to provide several candidate items in prompts for the LLM, e.g., the items recalled by traditional recommendation models (Lin et al., 2023b; Yue et al., 2023). Due to the input length limitation of LLMs, it is not possible to include all items in the prompts. In addition to selecting suitable candidate sets, there are also methods that directly generate candidate items by LLMs, utilizing strategies such as output probability distribution (Yue et al., 2023) and vector quantization (Zheng et al., 2023) for item indexing and grounding. Section 5.3 will focus on the construction strategies of candidate items, including selection and generation.
@@
$\bullet$ Prompting strategies. Despite the impressive capabilities of LLMs, they tend to exhibit unsatisfactory performance in providing personalized recommendations (Hou et al., 2023; Dai et al., 2023a; Liu et al., 2023c; Kang et al., 2023). The reason may stem from the significant semantic gap between the general knowledge encoded in LLMs and the domain-specific behavioral pattern and item catalogs of recommender systems (Zhang et al., 2023b; Li et al., 2023a). To specialize LLMs to recommender systems, we summarize and propose several prompting strategies specialized for recommendation tasks. Details will be discussed in Section 5.4.
Table 2. Instantiation of existing work for ProLLM4Rec.
@@
LLMs Task description User interest Candidate items Prompting strategies Related Work
Not tuning setting
@@
CTR predictions, rating, re-ranking recent and relevant items (with attributes), user profile pointwise, pairwise, listwise item(s) chain of thoughts, in-context learning, role prompting (Di Palma et al., 2023; Kang et al., 2023; Dai et al., 2023a; Shu et al., 2023; Hou et al., 2023; Wang and Lim, 2023; Wang et al., 2023c; Liu et al., 2023c; Yao et al., 2023; Zhiyuli et al., 2023; Liu et al., 2023d; Li et al., 2023k; Wang et al., 2023f; Liu et al., 2023e; Ma et al., 2023)
@@
ChatGPT, GPT-4 conversational recommender systems user explicit interest, interactive feedback recalled from traditional models role prompting (Gao et al., 2023; Friedman et al., 2023; He et al., 2023; Jin et al., 2023; Zhu et al., 2023; Wang et al., 2023h; Huang et al., 2023)
@@
generative recommendation recent items (with attributes), user profile (not provided, generation methods) basic prompts (Liu et al., 2023a; Wang et al., 2023e)
(Parameter-efficient) Fine-tuning setting
@@
LLaMA, LLaMA2, Vicuna, ChatGPT CTR predictions, rating, re-ranking recent and relevant items (with attributes), user profile, collaborative embedding pointwise, listwise item(s) chain of thoughts, role prompting, soft prompting (Bao et al., 2023b; Zhang et al., 2023b; Kang et al., 2023; Yang et al., 2023a; Lin et al., 2023b; Yue et al., 2023; Liao et al., 2023; Wang et al., 2023b; Di Palma et al., 2023; Li et al., 2023h; Fu et al., 2023; Shi et al., 2023)
@@
LLaMA, LLaMA2, BART, GPT recall, retrieving recent and relevant items (with attributes) (not provided, item grounding methods) basic prompts (Lin et al., 2023c; Bao et al., 2023a; Zheng et al., 2023; Li et al., 2023b; Petrov and Macdonald, 2023; Li et al., 2023m; Ji et al., 2023; Zhang et al., 2021)
Instruction tuning setting
@@
(Flan-)T5, LLaMA2 rating, ranking, retrieving, explanation, summarization recently interacted items, user profile, short-term intentions pointwise, pairwise, listwise item(s) basic prompts (Geng et al., 2022; Zhang et al., 2023f; Li et al., 2023j; Chu et al., 2023; Qiu et al., 2023; Luo et al., 2023b)
3.2. Instantiation of ProLLM4Rec
@@
By combining the key elements mentioned above, we can instantiate various types of recommender systems in our framework with the following five steps. Specifically, (1) we can employ LLMs with varying levels of public availability, different tuning strategies, model architectures, parameter scales, and context lengths. (2) We can define a range of task description, such as retrieving, rating, recalling, and ranking. (3) Regarding user interest modeling, we can employ different types of interest, representation forms, and modeling methods. (4) When collecting candidate items, we take into account their different representation types, sources, and grounding methods. (5) Additionally, we can introduce several well-designed prompting strategies to effectively guide the recommendation capabilities of LLMs. To demonstrate the compatibility and versatility of our framework, we summarize previous work on LLM-powered recommender systems in Table 2 based on various settings in our framework ProLLM4Rec.
@@ 3.3. Experimental Settings
In this section, we introduce the overall settings of the following experiments. We first describe the basic information of the two public datasets, and then present the configurations and implementations to conduct our study in detail.
@@ 3.3.1. Datasets
The domain characteristics of movies and books are closer to the general knowledge of LLMs, which facilitates the further analysis. Considering the scale, popularity and side information of public datasets, we select two representative datasets to conduct our study, i.e., MovieLens-1M (Harper, 2015) and Amazon Books (2018) (Ni et al., 2019) as follows.
$\bullet$ MovieLens-1M (Harper, 2015) is one of the most widely used benchmark datasets in the field of recommender systems, covering movie ratings and attributes on the website movielens.org. We use the one million version from the MovieLens datasets.
$\bullet$ Amazon Books (2018) (Ni et al., 2019) is an updated version of the Amazon review dataset released in 2014. At first, Amazon only operated online book sales business, so the data in the book field is the most abundant. To improve the data quality, we filter out inactive users and unpopular products, and remove dirty data without necessary attributes.
In our research, we are concerned about how LLMs can fully utilize the domain knowledge to make recommendations, and use the title of items as the input for prompts. However, titles are not enough to describe items, and there are deviations between the text of the title and the content of the item itself (e.g., the movie Twelve Monkeys). Therefore, we further investigate the benefits of detailed item descriptions on the recommendation effect. As shown in Table 3, there are no item descriptions in the original dataset of MovieLens, only the release year, title, and genre. To enrich the movie dataset, we use the general knowledge of ChatGPT  [^1] The URL of ChatGPT API: https://chat.openai.com/. Note that there are multiple versions of the ChatGPT API, and OpenAI has released interfaces on March 1 and June 13, 2023, respectively. In the absence of clear annotations, the ChatGPT used in this article is “gpt3.5-turbo-4k-0613”. to generate text descriptions for movies.
Table 3. Statistics of two public datasets for ProLLM4Rec.
@@
Amazon-Books 13,469 12,984 1,142,940 0.6536% title, categories, brand, price, description

### v2

@@ 2.3. Differences with Existing Work
Previous surveys of LLMs for recommender system usually categorized existing work with a classification standard and introduced studies. Lin et al.  (Lin et al., 2023a) categorized existing work into the targets and the methods of adapting LLMs to recommendation tasks. Wu et al.  (Wu et al., 2024b) mainly focused on the form of information between LLMs and recommender systems. Zhao et al.  (Zhao et al., 2024) summarized the framework into four sections, involving deep representation learning, pre-training, fine-tuning and prompting of LLMs. Li et al.  (Li et al., 2024d) concentrated on LLMs for generative recommendation, and summarized the different recommendation tasks that LLMs can serve as in detail. Chen et al.  (Chen et al., 2024b) analyzed challenges and opportunities of LLMs in the field of personalization, and Li et al.  (Li et al., 2024b) systematically classified existing work on generative retrieval and generative recommendation. These surveys mainly concentrated on demonstrating related work and summarizing the advantages and limitations, while they lack performance analysis supported by experimental results. Although existing work like LLMRec (Liu et al., 2023d) and LLM-REC (Tan et al., 2024) has evaluated the recommendation performance of LLMs on multiple tasks, their analysis of LLMs selection and prompt design is not deep enough. Compared to previous work, we concentrate on the performance of LLMs leveraged as recommender systems, and provide a systematic empirical analysis on LLM-based recommendations by devising a general framework LLM-RS. We mainly focus on two aspects, i.e., LLMs and task-specific prompts, providing definitions and solutions from both conceptual and methodological perspectives. Furthermore, we conduct experiments to discover new findings and validate results previously discussed in existing research, serving as an inspiration for future research efforts.
@@ 3. GENERAL FRAMEWORK AND OVERALL SETTINGS
Figure 1. The overall framework of our proposed LLM-RS. In our framework, LLMs are leveraged as recommender systems in four ways: prompting without tuning, full-model fine-tuning, parameter-efficient fine-tuning and instruction tuning; Prompt engineering consists of four components: task description, user interest modeling, candidate items construction and prompting strategies. LLMs act as recommender systems through task-specific prompts, while LLMs provide response and feedback to optimize prompts.
[…]
The framework consists of five components: (1) LLMs Leveraged as Recommender Systems; (2) Task Description; (3) User Interest; (4) Candidate Items and (5) Prompting Strategies. (2) - (5) constitute Task-specific Prompts.
In this section, we present our proposed LLM-RS with two important components, namely LLMs and prompts. Generally speaking, the rich world knowledge and general capabilities of LLMs demonstrate the potential to develop LLM-based recommender systems. Nevertheless, it is essential to introduce appropriate prompting engineering to provide domain-specific knowledge on recommendation tasks for LLMs to serve as personalized recommenders. In the following, we first describe our general framework by defining several key elements (Section 3.1). Then, we explain how our framework can be generalized to various tasks by framework instantiation (Section 3.2). Finally, we introduce our overall experimental settings for further analysis (Section 3.3).
@@ 3.1. Overview of the Approach
When it comes to leveraging LLMs as recommender systems, the first step involves choosing suitable LLMs tailored to specific scenarios, with an emphasis on their distinct capabilities. In our framework, LLMs are leveraged as recommender systems in four ways: prompting without tuning, full-model fine-tuning, parameter-efficient fine-tuning and instruction tuning. Subsequently, it is important to conduct prompt engineering to instruct LLMs to perform effective recommendations. Typically, in the realm of recommender systems using LLMs, despite the task settings in existing studies vary a lot, the overall prompting format remains relatively consistent with minor variations (Hou et al., 2024; Gao et al., 2023; Dai et al., 2023a; Wang et al., 2024c). Therefore, to unify existing prompting approaches for different recommendation purposes, as illustrated in Fig. 1, we establish a general framework for prompt engineering in LLM-RS, consisting of four key factors in the prompts. Firstly, task description is required to clearly express the specific goal of recommendation tasks. Secondly, prompts need to be carefully designed, to express user interest and meanwhile enable LLMs to provide personalized recommendations with both world and domain knowledge. Thirdly, the purpose of a recommender system is to provide users with candidate items, and potential candidate items should be constructed for LLMs to facilitate the recommendation with the understanding of domain-specific item information. Furthermore, special prompting strategies can be further employed to enhance the specialized recommendation capabilities of LLMs. In what follows, we will thoroughly examine the impact of each factor on the recommendation performance.
@@ 3.1.1. Key Elements in Our Framework
To carry out the experiments, we first describe the key elements in LLM-RS, to clarify their definitions and scope in this work. Specially, we introduce the following five elements for our framework:
@@
$\bullet$ Large language models (LLMs). As proposed in previous research (Hou et al., 2024; Zhang et al., 2023b; Li et al., 2023a; Bao et al., 2023a), there exists a large gap between general language modeling and personalized user behavioral modeling, making it non-trivial to utilize LLMs in recommender systems. In this work, we investigate the efficacy of LLMs from perspectives of public availability, tuning strategies, model architecture, parameter scale, and context length, aiming to gain insights into the selection of appropriate LLMs for performing recommendation tasks. Observations and discussions on the use of LLMs are presented in Section 4.
@@
$\bullet$ Task description. To adapt LLMs to the scenario of recommendation, it is necessary to clearly express the context and target of recommender systems for LLMs in the prompt, i.e., task description. With different prompt descriptions of tasks, LLMs can be leveraged to various recommendation scenarios and tasks such as click-through rate predictions (Kang et al., 2023; Bao et al., 2023b, a), sequential recommendation (Hou et al., 2024; Dai et al., 2023a; Zheng et al., 2024) and conversational recommender systems (He et al., 2023; Gao et al., 2023; Wang et al., 2023f). In addition, the well-designed task description also helps to improve the fairness (Wang et al., 2024d; Zhang et al., 2023a) and diversity (Carraro and Bridge, 2024; Deldjoo, 2024) of recommendation results.
@@
$\bullet$ User interest modeling. The modeling of user interest is the key to recommendation tasks (Rendle et al., 2009; He et al., 2017). From a temporal perspective, user interest can be categorized into short-term instant interest and long-term stable preferences, which collectively constitute the user profile (Tan et al., 2024; Liu et al., 2023d). When leveraging LLMs as recommender systems, users are generally expressed in natural language text (Hou et al., 2024; Bao et al., 2023b), which is different from traditional approaches capturing user preference from ID-based behavior sequences (Kang and McAuley, 2018; Wu et al., 2021a). Moreover, learnable tokens (Wang et al., 2024a; Kim et al., 2024; Shao et al., 2024) and distilled soft prompts (Ramos et al., 2024; Li et al., 2023g) for user interest modeling in LLMs have attracted attention of researchers. In this paper, we mainly consider the reflected user interest based on his or her interaction behaviors with interacted items. Especially, as detailed in Section 5.2, we employ item description texts, user profiles, and historical interactions between users and items to reveal the underlying user interest using natural languages (Wang et al., 2024c; Lin et al., 2024a; Shu et al., 2023; Yao et al., 2023).
@@
$\bullet$ Candidate items construction. The purpose of recommender systems is to provide users with items to choose from, so candidate items construction is a crucial step in our framework (Hou et al., 2024; Zhang et al., 2023e; Dai et al., 2023a). A simple approach is to provide several candidate items in prompts, e.g., the items recalled by traditional recommendation models (Lin et al., 2024a; Yue et al., 2023). Due to the input length limitation of LLMs, it is not possible to include all items in the prompts. Besides selecting candidate sets, there are also methods that directly generate candidate items by LLMs, utilizing strategies such as output probability distribution (Yue et al., 2023) and vector quantization (Zheng et al., 2024) for item indexing and grounding. Section 5.3 will focus on the construction strategies of candidate items, including selection, grounding and generation.
@@
$\bullet$ Prompting strategies. Despite the impressive capabilities of LLMs, they tend to exhibit unsatisfactory performance in providing personalized recommendations (Hou et al., 2024; Dai et al., 2023a; Liu et al., 2023c; Kang et al., 2023). The reason may stem from the significant semantic gap between the general knowledge encoded in LLMs and the domain-specific behavioral pattern and item catalogs of recommender systems (Zhang et al., 2023b; Li et al., 2023a). To specialize LLMs to recommender systems, we summarize and propose several prompting strategies specialized for recommendation tasks. Details will be discussed in Section 5.4.
Table 2. Instantiation of existing work for LLM-RS. We summarize three paradigms of leveraging LLMs as recommender systems within the framework of our proposed LLM-RS: (1) prompting without tuning, (2) (parameter-efficient) fine-tuning, and (3) instruction tuning. For each paradigm, we provide a comprehensive overview of related work and its instantiation in our framework.
@@
LLMs Task Descriptions User Interest Candidates Prompt Strategies Related Work
Prompting without Tuning
@@
CTR predictions, rating, re-ranking recent and relevant items (with attributes), user profile pointwise, pairwise, listwise item(s) chain-of-thoughts, in-context learning, self-inspiring, role prompting (Di Palma et al., 2023; Kang et al., 2023; Dai et al., 2023a; Shu et al., 2023; Hou et al., 2024; Wang and Lim, 2023; Wang et al., 2024c; Liu et al., 2023c; Yao et al., 2023; Li et al., 2023b; Liu et al., 2023d; Li et al., 2023j; Wang et al., 2023d; Liu et al., 2024d; Ma et al., 2023; Carraro and Bridge, 2024; Spurlock et al., 2024; Liu et al., 2024b; Sanner et al., 2023b)
@@
ChatGPT, GPT-4, GPT-4o conversational recommender systems user explicit interest, interactive feedback recalled from traditional models role prompting (Gao et al., 2023; Friedman et al., 2023; He et al., 2023; Jin et al., 2023; Wang et al., 2023f; Huang et al., 2023)
@@
generative recommendation recent items (with attributes), user profile (not provided, generation methods) basic prompts (Liu et al., 2023a; Wang et al., 2023c)
(Parameter-efficient) Fine-tuning
@@
LLaMA, Vicuna, Flan-T5, BART, GPT, PaLM CTR predictions, rating, re-ranking recent and relevant items (with attributes), user profile, collaborative embedding pointwise, listwise item(s) chain of thoughts, role prompting, soft prompting (Bao et al., 2023b; Zhang et al., 2023b; Kang et al., 2023; Yang et al., 2023a; Lin et al., 2024a; Yue et al., 2023; Liao et al., 2023; Wang et al., 2023b; Di Palma et al., 2023; Li et al., 2023h; Fu et al., 2023; Shi et al., 2024; Qin et al., 2024; Wu et al., 2024a; Mao et al., 2023)
@@
recall, retrieving tasks, conversational recommender systems recent and relevant items (with attributes) (not provided, item grounding methods) basic prompts (Lin et al., 2023b; Bao et al., 2023a; Zheng et al., 2024; Li et al., 2023c; Petrov and Macdonald, 2023; Li et al., 2023k; Ji et al., 2023; Zhang et al., 2021; Zhu et al., 2024; Feng et al., 2023; Yang et al., 2022; Lin et al., 2024b; Rajput et al., 2023; Chen et al., 2024a; Wang et al., 2024a; Yang et al., 2024; Li et al., 2024e; Paischer et al., 2024; Yin et al., 2024; Penha et al., 2024; Li et al., 2024c; Ding et al., 2024; Kim et al., 2024; Shao et al., 2024; Liu et al., 2024c)
Instruction Tuning
@@
(Flan-)T5, LLaMA rating, ranking, explanation generation, review summarization recent items, user profile, short-term intentions pointwise, pairwise, listwise item(s) basic prompts (Geng et al., 2022; Zhang et al., 2023e; Li et al., 2023i; Chu et al., 2023; Qiu et al., 2023; Luo et al., 2023; Li et al., 2024a, 2023g; Hua et al., 2023; Geng et al., 2023; Wei et al., 2024a; Ramos et al., 2024; Tan et al., 2024; Wang et al., 2024b)
3.2. Instantiation of LLM-RS
@@
By combining the key elements mentioned above, we can instantiate various types of recommender systems in our framework with the following five steps. Specifically, (1) we can employ LLMs with varying levels of public availability, different tuning strategies, model architectures, parameter scales, and context lengths. (2) We can define a range of task description, such as retrieving, rating, recalling, and ranking. (3) Regarding user interest modeling, we can employ different types of interest, representation forms, and modeling methods. (4) When collecting candidate items, we take into account their different representation types, sources, and grounding methods. (5) Additionally, we can introduce several well-designed prompting strategies to effectively guide the recommendation capabilities of LLMs. To demonstrate the compatibility and versatility of our framework, we summarize previous work on LLM-based recommender systems in Table 2 based on various settings in our framework LLM-RS.
@@ 3.3. Experimental Settings
In this section, we introduce the overall settings of the following experiments. We first describe the basic information of datasets, and then present the configurations and implementations.
@@ 3.3.1. Datasets
The domain characteristics of movies and books are closer to the general knowledge of LLMs, which facilitates the further analysis. Considering the scale, popularity and side information of public datasets, we select two representative datasets to conduct our study, i.e., MovieLens-1M (Harper and Konstan, 2016) and Amazon Books (2018) (Ni et al., 2019). Next, we present the details of the two selected datasets:
$\bullet$ MovieLens-1M (Harper and Konstan, 2016) is one of the most widely used benchmark datasets in the field of recommender systems, covering movie ratings and attributes on the website movielens.org. We use the one million version from the MovieLens datasets, and it contains 1,000,209 ratings from 6,040 users on 3,706 movies. It is collected by the GroupLens research group.
$\bullet$ Amazon Books (2018) (Ni et al., 2019) is an updated version of the Amazon review dataset. Amazon review datasets include reviews and product metadata collected from Amazon, one of the largest online e-commerce company in the United States. At first, Amazon only operated online book sales business, so the data in the book field is the most abundant. To improve the data quality, we filter out inactive users and unpopular products, and remove data without necessary attributes.
In our research, we are concerned about how LLMs can fully utilize the domain knowledge to make recommendations, and use the title of items as the input for prompts. However, titles are not enough to describe items, and there are deviations between the text of the title and the content of the item itself (e.g., the movie Twelve Monkeys). Therefore, we further investigate the benefits of detailed item descriptions on the recommendation effect. As shown in Table 3, there are no item descriptions in the original dataset of MovieLens, only the release year, title, and genre. If LLMs do not include knowledge about a movie, only the release year, title and genre may lead to ambiguity and hallucination. To enrich the movie dataset, we use the general knowledge of ChatGPT  [^1] The URL of ChatGPT Application Programming Interface (API): https://chat.openai.com/. Note that there are multiple versions of the ChatGPT API. In the absence of clear annotations, the ChatGPT used in this article is “gpt3.5-turbo-4k-0613”. to generate text descriptions for movies. We mainly use the interaction information between users and items in the recommendation dataset, as well as the title and description of items.
Table 3. Statistics of two public datasets for LLM-RS. Since the original MovieLens-1M dataset does not provide description information, we used ChatGPT to generate text descriptions based on item attributes.
@@
Amazon Books 13,469 12,984 1,142,940 0.6536% title, categories, brand, price, description